One number is 4 greater than another number. When three times the smaller number is added to twice the larger one, the result in 43. Find the numbers.
What is the formula for this question?

Answer by stanbon(75887)

One number is 4 greater than another number. When three times the smaller number is added to twice the larger one, the result in 43. Find the numbers.
---
If x = y + 5, y must be the smaller number.
3y + 2x = 43
-----
Substitute for "x" and solve for "y":
3y + 2(y+5) = 43
3y + 2y + 10 = 43
5y = 33
y = 33/6
-------
Solve for "x":
x = y + 5
x = (33/6) + (30/6)
x = 63/6
=====================
Cheers,
Stan H.
=====================